# Joint Learning of Energy-based Models and their Partition Function

### v1

@@ 4.1 Multilabel classification
$[{\bm{y}}^{\star}_{g}({\bm{x}})]_{j}=\begin{cases}1&\mbox{ if }\theta_{j}\geq 0\\ 0&\mbox{ if }\theta_{j}<0\end{cases}\quad\forall j\in[k].$ (45)
@@ Results.
We evaluate our models on classical multilabel classification datasets. The hyperparameters (learning rate and regularization) are optimized on a validation set. Once the best hyperparameters are found, we refit the model on the combined training and validation sets and evaluate it on the test set. Our results in Table 3 show that the logistic and sparsemax losses trained with our approach work better than the generalized Fenchel-Young loss in various configurations. We also present learning curves in Figure 1. We numerically validate that our approach converges to true MLE as more ${\bm{y}}^{\prime}$ samples are used. We also observe a regularization phenomenon: using fewer ${\bm{y}}^{\prime}$ samples often leads to better $f_{1}$ scores than exact MLE. These findings are confirmed on all other datasets we considered; see Figure 3.
@@ 4.2 Label ranking
A permutation can be represented as a vector. In this case, $\operatorname*{conv}(\mathcal{Y})=\mathcal{P}$, where $\mathcal{P}$ is the permutahedron (Bowman, 1972; Ziegler, 2012; Blondel et al., 2020b). We set the logits to ${\bm{\theta}}=h({\bm{x}})\in\mathbb{R}^{k}$. The argmax
${\bm{y}}^{\star}_{g}({\bm{x}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{Y}}\Phi({\bm{\theta}},{\bm{y}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{P}}\langle{\bm{\theta}},{\bm{y}}\rangle$ (51)
can be solved optimally using an argsort (Sander et al., 2023, Proposition 1). However, $\mathrm{LSE}_{g}$ and $\mu_{g}$ are known to be #P-complete to compute, making classical MLE intractable.
@@
A permutation can also be represented as a permutation matrix. In this case, $\operatorname*{conv}(\mathcal{Y})=\mathcal{B}$, where $\mathcal{B}$ is the Birkhoff polytope, the set of doubly-stochastic matrices. The logits are ${\bm{\theta}}=h({\bm{x}})\in\mathbb{R}^{k\times k}$. The argmax
${\bm{y}}^{\star}_{g}({\bm{x}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{Y}}\Phi({\bm{\theta}},{\bm{y}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{B}}\langle{\bm{\theta}},{\bm{y}}\rangle$ (52)
can be solved optimally using the Hungarian algorithm (Kuhn, 1955). However, $\mathrm{LSE}_{g}$ and $\mu_{g}$ are again #P-complete to compute, making classical MLE intractable.
@@ Acknowledgements
We thank Alexandre Ramé for feedback on this paper.
@@
Author contributions
  • Michaël E. Sander: implementation, experiments, proofs, writing.
  • Vincent Roulet: proofs, writing, code reviews.
  • Tianlin Liu: writing, code reviews.
  • Mathieu Blondel: initial project idea, implementation, experiments, proofs, writing.
@@ Appendix A Experimental details and additional results
Figure 3: Convergence of the proposed approach as a function of the number of prior samples ${\bm{y}}^{\prime}$ drawn. To be able to compute the exact MLE objective (27), we use the unary multilabel model (Section 4.1) as the test bed. Note that the loss and gradient in plots (a) and (b) are computed using (27) even for our method. We make two key observations: i) our approach converges to exact MLE as predicted by our theory, ii) the number of ${\bm{y}}^{\prime}$ samples can have a regularization effect on the test $f_{1}$-score.
@@
Additional experimental results are shown in Tables 3, 4 and 5.
Loss Model $g$ yeast scene birds emotions cal500
logistic Linear (unary) 63.41 70.78 45.49 61.63 40.19
MLP (unary) 65.04 75.35 46.48 65.33 44.91
ResNet (unary) 65.03 75.64 36.93 67.45 45.23
Linear (pairwise) 63.40 70.72 45.49 61.63 40.23
MLP (pairwise) 65.11 75.22 46.08 64.84 45.00
ResNet (pairwise) 64.84 75.68 41.34 64.26 44.61
sparsemax Linear (unary) 62.92 70.43 45.12 63.14 45.18
MLP (unary) 63.22 75.83 42.42 66.86 46.39
ResNet (unary) 62.27 75.44 28.63 44.53 39.11
Linear (pairwise) 62.91 69.89 45.12 63.14 45.18
MLP (pairwise) 63.39 74.94 40.68 66.28 46.39
ResNet (pairwise) 62.32 75.63 27.85 64.58 30.93
GFY Linear (unary) 61.41 70.26 44.13 61.70 36.85
MLP (unary) 61.31 70.84 40.73 61.85 37.35
Linear (pairwise) 61.62 70.42 42.28 63.05 35.12
@@ B.1 Loss functions using other forms of supervision
where $\mathrm{sigmoid}(u)\coloneqq 1/(1+\exp(u))$ is the logistic sigmoid. The loss associated with the triplet $({\bm{x}},{\bm{y}}_{+},{\bm{y}}_{-})$ is the pairwise logistic loss, $-\log p_{g}({\bm{y}}_{+}\succ{\bm{y}}_{-}|{\bm{x}})$. A major advantage of this approach is that there is no log-partition involved. The DPO loss (Rafailov et al., 2024) can be thought as a generalization of this pairwise logistic loss to non-uniform prior distributions $q$. However, it learns a probabilistic model of ${\bm{y}}_{1}\succ{\bm{y}}_{2}$ given ${\bm{x}}$, not of ${\bm{y}}$ given ${\bm{x}}$.
@@ C.1 Lemmas
Proof. This is a corollary of Lemma C.1. $\square$
@@
Proposition 3.1 is a corollary of Proposition 39 with the choice
@@ C.3 Proof of Proposition 28 (optimality in the empirical risk setting)
where $\delta_{\bm{y}}\in\mathcal{P}_{1}(\mathcal{Y})$ is a delta distribution such that $\delta_{\bm{y}}({\bm{y}}^{\prime})=1$ for ${\bm{y}}^{\prime}={\bm{y}}$ and 00 for ${\bm{y}}^{\prime}\neq{\bm{y}}$. Choosing (67) as in the proof of Proposition 3.1 gives the final result.
C.4 Proof of Proposition 32 (min-max formulation)
@@
$\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathcal{L}(g)$ $=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathbb{E}_{\bm{x}}\left[\Omega^{*}(g_{\bm{x}})-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]$ (77)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathbb{E}_{\bm{x}}\left[\left(\max_{p_{\bm{x}}\in\mathcal{P}_{1}(\mathcal{Y})}\langle g_{\bm{x}},p_{\bm{x}}\rangle-\Omega(p_{\bm{x}})\right)-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]$ (78)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\max_{p\in\mathcal{P}_{1}(\mathcal{Y}|\mathcal{X})}\mathbb{E}_{\bm{x}}\left[(\langle g_{\bm{x}},p_{\bm{x}}\rangle-\Omega(p_{\bm{x}}))-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]$ (79)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\max_{p\in\mathcal{P}_{1}(\mathcal{Y}|\mathcal{X})}\mathbb{E}_{\bm{x}}\left[\left(\mathbb{E}_{{\bm{y}}\sim p_{\bm{x}}}[g({\bm{x}},{\bm{y}})]-\Omega(p_{\bm{x}})\right)-\mathbb{E}_{{\bm{y}}\sim\rho_{\bm{x}}}[g({\bm{x}},{\bm{y}})]\right].$ (80)
C.5 Proof of Proposition 39 (optimality in the general Fenchel-Young loss and expected risk setting)
@@
We can then analyze the following function on $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]\times\mathcal{X}$:
@@
First, we need to ensure that we can compute derivatives of this function with respect to its first variable $\tau$ in $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. We have that $f^{*}(\max\{g({\bm{x}},{\bm{y}})-\tau,f^{\prime}(0)\})=(f+\iota_{\mathbb{R}_{+}})^{*}(g({\bm{x}},{\bm{y}})-\tau)$ and $\operatorname*{dom}(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}=\operatorname*{dom}f_{*}^{\prime}\cup(-\infty,f^{\prime}(0)]$. Therefore, if $f^{\prime}(0)>-\infty$, the domain of $(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$ is unbounded below. Otherwise, if $f^{\prime}(0)=-\infty$, since $\mathrm{Im}f^{\prime}\subseteq\operatorname*{dom}f_{*}^{\prime}$, the domain of $f_{*}^{\prime}$, and so of $(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$, are unbounded below. Denoting then $\alpha=\sup\operatorname*{dom}(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$, since $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\min}({\bm{x}})=f^{\prime}(\frac{1}{\min_{{\bm{y}}\in\mathcal{Y}}q({{\bm{y}}|{\bm{x}}})})\in\operatorname*{dom}f_{*}^{\prime}$, we have $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\min}({\bm{x}})<\alpha$ and therefore
@@
=$\stackrel{{\scriptstyle(ii)}}{{=}}1-\frac{q({\bm{y}}^{\star}|{\bm{x}})}{\min_{y\in\mathcal{Y}}(q({\bm{y}},{\bm{x}}))}\leq 0,$
@@
Finally, we show that $h^{\prime}$ is increasing on $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. When $\tau\leq\tau_{\max}({\bm{x}})$, one has $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau\geq\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\max}(x)\geq f^{\prime}(0).$
Since $\theta\mapsto f_{*}^{\prime}(\max\{\theta,f^{\prime}(0)\})$ is increasing on $\operatorname*{dom}f^{*}\setminus(-\infty,f^{\prime}(0)]$ we have that $\tau\mapsto-f_{*}^{\prime}(\max\{(g({\bm{x}},{\bm{y}}^{\star})-\tau,f^{\prime}(0)\})$ is increasing on $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$ and therefore $\tau\mapsto h^{\prime}(\tau,{\bm{x}})=1-\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})f_{*}^{\prime}(\max\{g({\bm{x}},{\bm{y}})-\tau,f^{\prime}(0)\})$ is increasing on $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$.
Overall $h$ is well defined and strictly convex on $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$ such that $h^{\prime}(\tau_{\min}({\bm{x}}))\leq 0$ and $h^{\prime}(\tau_{\max}({\bm{x}}))\geq 0$. Hence, since $h$ is convex, we have
@@
and the unique minimizer $\tau^{\star}({\bm{x}})$ can then be found by solving the first order optimality condition $h^{\prime}(\tau,{\bm{x}})=0$ in $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. We can therefore define a function $\tau^{\star}$ mapping each ${\bm{x}}$ to the unique root of $h^{\prime}(\tau,{\bm{x}})$ in $[\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$.

### v3

@@ 4.1 Multilabel classification
${\bm{y}}^{\star}_{g}({\bm{x}})]_{j}=\begin{cases}1&\mbox{ if }\theta_{j}\geq 0\\ 0&\mbox{ if }\theta_{j}<0\end{cases}\quad\forall j\in[k].$ (42)
@@
Logistic (min-min) $\mathcal{P}$ Linear 85.14 80.47 57.78 79.74 50.82 91.36
$\mathcal{P}$ MLP 88.69 84.39 97.78 86.27 74.26 90.12
$\mathcal{B}$ Linear 87.71 81.29 97.78 83.33 56.33 95.06
$\mathcal{B}$ MLP 90.80 85.43 98.52 86.41 65.20 91.98
Sparsemax (min-min) $\mathcal{P}$ Linear 84.88 79.84 52.59 78.76 50.25 93.83
$\mathcal{P}$ MLP 86.52 85.94 99.26 85.75 68.34 91.36
$\mathcal{B}$ Linear 89.09 79.43 97.78 82.81 56.24 90.12
$\mathcal{B}$ MLP 87.84 80.47 98.52 87.19 55.43 85.80
Logistic (MCMC) $\mathcal{P}$ Linear 84.29 80.67 55.56 79.15 49.97 91.36
$\mathcal{P}$ MLP 87.90 75.50 97.04 86.34 57.90 91.98
$\mathcal{B}$ Linear 88.10 79.33 97.78 82.22 55.99 96.91
$\mathcal{B}$ MLP 89.68 80.16 97.78 85.88 48.90 91.98
Logistic (min-max) $\mathcal{P}$ Linear 61.01 79.22 61.48 70.20 45.68 65.43
$\mathcal{P}$ MLP 70.02 67.03 55.56 68.56 40.22 70.99
$\mathcal{B}$ Linear 82.58 78.40 61.48 74.90 52.86 93.21
$\mathcal{B}$ MLP 78.50 71.68 42.96 68.76 48.99 83.33
@@ Results.
We evaluate our models on classical multilabel classification datasets. The hyperparameters (learning rate and regularization) are optimized on a validation set. Once the best hyperparameters are found, we refit the model on the combined training and validation sets and evaluate it on the test set. Our results in Table 1 show that the logistic and sparsemax losses trained with our approach work better than the generalized Fenchel-Young loss as well as min-max and MCMC sampling approaches in various configurations. For the min-max approach, we use optimistic ADAM as solver, an MLP as generator and we use REINFORCE (score function estimator) for gradient estimation. For MCMC sampling, we use standard Metropolis–Hastings algorithm with uniform proposal distribution. We also present learning curves in Figure 1. We numerically validate that our approach converges to true MLE as more $\bm{y}$$^{\prime}$ samples are used. We also observe a regularization phenomenon: using fewer $\bm{y}$$^{\prime}$ samples leads to better $f_{1}$ scores than exact MLE. These findings are confirmed on other datasets (Figure 3).
@@ 4.2 Label ranking
A permutation can be represented as a vector. In this case, $\operatorname*{conv}(\mathcal{Y})=\mathcal{P}$, where $\mathcal{P}$ is the permutahedron (Bowman, 1972; Ziegler, 2012; Blondel et al., 2020b). We set the logits to ${\bm{\theta}}=h({\bm{x}})\in\mathbb{R}^{k}$. The argmax ${\bm{y}}^{\star}_{g}({\bm{x}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{Y}}\Phi({\bm{\theta}},{\bm{y}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{P}}\langle{\bm{\theta}},{\bm{y}}\rangle$ can be solved optimally using an argsort (Sander et al., 2023, Proposition 1). However, $\mathrm{LSE}_{g}$ and ${\bm{\mu}}_{g}$ are known to be #P-complete to compute, making classical MLE intractable.
@@
A permutation can also be represented as a permutation matrix. In this case, $\operatorname*{conv}(\mathcal{Y})=\mathcal{B}$, where $\mathcal{B}$ is the Birkhoff polytope, the set of doubly-stochastic matrices. The logits are ${\bm{\theta}}=h({\bm{x}})\in\mathbb{R}^{k\times k}$. The argmax ${\bm{y}}^{\star}_{g}({\bm{x}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{Y}}\Phi({\bm{\theta}},{\bm{y}})=\operatorname*{argmax}_{{\bm{y}}\in\mathcal{B}}\langle{\bm{\theta}},{\bm{y}}\rangle$ can be solved optimally using the Hungarian algorithm (Kuhn, 1955). However, $\mathrm{LSE}_{g}$ and $\mu_{g}$ are again #P-complete to compute, making classical MLE intractable.
@@ 5 Conclusion and future work
In this paper, we evaluated our proposed method primarily in the prediction setting (finding the mode). Our paper does not directly address the challenge of sampling from $p({\bm{y}}|{\bm{x}})$, which is central to evaluating performance in a probabilistic or generative setting. We leave to future work whether the learned log-partition function can be used to design better samplers.
@@ Acknowledgements
We thank Alexandre Ramé as well as the anonymous reviewers for constructive feedbacks on this paper.
@@ Appendix A Experimental details and additional results
Figure 3: Convergence of the proposed approach as a function of the number of prior samples $\bm{y}$$^{\prime}$ drawn. To be able to compute the exact MLE objective (27), we use the unary multilabel model (Section 4.1) as the test bed. Note that the loss and gradient in the left and center columns are computed using (27) even for our method. We make two key observations: i) our approach converges to exact MLE as predicted by our theory, ii) the number of $\bm{y}$$^{\prime}$ samples can have a regularization effect on the test $f_{1}$-score.
@@
Additional experimental results are shown in Tables 4 and 5.
@@ B.1 Loss functions using other forms of supervision
where $\mathrm{sigmoid}(u)\coloneqq 1/(1+\exp(u))$ is the logistic sigmoid. The loss associated with the triplet $({\bm{x}},{\bm{y}}_{+},{\bm{y}}_{-})$ is the pairwise logistic loss, $\log p_{g}({\bm{y}}_{+}\succ{\bm{y}}_{-}|{\bm{x}})$. A major advantage of this approach is that there is no log-partition involved. The DPO loss (Rafailov et al., 2024) can be thought as a generalization of this pairwise logistic loss to non-uniform prior distributions $q$. However, it learns a probabilistic model of ${\bm{y}}_{1}\succ{\bm{y}}_{2}$ given $\bm{x}$}, not of $\bm{y}$} given $\bm{x}$}.
@@ C.1 Lemmas
Proof. This follows from classical conjugate calculus. $\square$
@@
Proposition 3.1 is a corollary of Proposition 3.4 with the choice
@@ C.3 Proof of Proposition 28 (optimality in the empirical risk setting)
where $\delta_{\bm{y}}\in\mathcal{P}_{1}(\mathcal{Y})$ is a delta distribution such that $\delta_{\bm{y}}({\bm{y}}^{\prime})=1$ for ${\bm{y}}^{\prime}={\bm{y}}$ and 0 for ${\bm{y}}^{\prime}\neq{\bm{y}}$. Choosing (62) as in the proof of Proposition 3.1 gives the final result.
C.4 Proof of Proposition 3.3 (min-max formulation)
@@
$\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathcal{L}(g)$ $=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathbb{E}_{\bm{x}}\left[\mathrm{LSE}_{g}(x)-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]+\mathrm{const}$ (72)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\mathbb{E}_{\bm{x}}\left[\left(\max_{p_{\bm{x}}\in\mathcal{P}_{1}(\mathcal{Y})}\langle g_{\bm{x}},p_{\bm{x}}\rangle-\mathrm{KL}(p_{\bm{x}},q_{\bm{x}})\right)-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]+\mathrm{const}$ (73)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\max_{p\in\mathcal{P}_{1}(\mathcal{Y}|\mathcal{X})}\mathbb{E}_{\bm{x}}\left[(\langle g_{\bm{x}},p_{\bm{x}}\rangle-\mathrm{KL}(p_{\bm{x}},q_{\bm{x}}))-\langle g_{\bm{x}},\rho_{\bm{x}}\rangle\right]+\mathrm{const}$ (74)
$=\min_{g\in\mathcal{F}(\mathcal{X}\times\mathcal{Y})}\max_{p\in\mathcal{P}_{1}(\mathcal{Y}|\mathcal{X})}\mathbb{E}_{\bm{x}}\left[\left(\mathbb{E}_{{\bm{y}}\sim p_{\bm{x}}}[g({\bm{x}},{\bm{y}})]-\mathrm{KL}(p_{\bm{x}},q_{\bm{x}})\right)-\mathbb{E}_{{\bm{y}}\sim\rho_{\bm{x}}}[g({\bm{x}},{\bm{y}})]\right]+\mathrm{const}.$ (75)
C.5 Proof of Proposition 3.4 (optimality in the general Fenchel-Young loss and expected risk setting)
@@
We can then analyze the following function on $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]\times\mathcal{X}$:
@@
First, we need to ensure that we can compute derivatives of this function with respect to its first variable $\tau$ in $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. We have that $f^{*}(\max\{g({\bm{x}},{\bm{y}})-\tau,f^{\prime}(0)\})=(f+\iota_{\mathbb{R}_{+}})^{*}(g({\bm{x}},{\bm{y}})-\tau)$ and $\operatorname*{dom}(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}=\operatorname*{dom}f_{*}^{\prime}\cup(-\infty,f^{\prime}(0)]$. Therefore, if $f^{\prime}(0)>$-$\infty$, the domain of $(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$ is unbounded below. Otherwise, if $f^{\prime}(0)=$-$\infty$, since $\mathrm{Im}f^{\prime}\subseteq\operatorname*{dom}f_{*}^{\prime}$, the domain of $f_{*}^{\prime}$, and so of $(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$, are unbounded below. Denoting then $\alpha=\sup\operatorname*{dom}(f+\iota_{\mathbb{R}_{+}})_{*}^{\prime}$, since $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\min}({\bm{x}})=f^{\prime}(\frac{1}{\min_{{\bm{y}}\in\mathcal{Y}}q({{\bm{y}}|{\bm{x}}})})\in\operatorname*{dom}f_{*}^{\prime}$, we have $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\min}({\bm{x}})<\alpha$ and therefore
@@
$\stackrel{{\scriptstyle(ii)}}{{=}}1-\frac{q({\bm{y}}^{\star}|{\bm{x}})}{\min_{y\in\mathcal{Y}}(q({\bm{y}},{\bm{x}}))}\leq 0,$
@@
Finally, we show that $h^{\prime}$ is increasing on $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. When $\tau\leq\tau_{\max}({\bm{x}})$, one has $\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau\geq\max_{{\bm{y}}\in\mathcal{Y}}g({\bm{x}},{\bm{y}})-\tau_{\max}(x)\geq f^{\prime}(0).$
Since $\theta\mapsto f_{*}^{\prime}(\max\{\theta,f^{\prime}(0)\})$ is increasing on $\operatorname*{dom}f^{*}\setminus(-\infty,f^{\prime}(0)]$ we have that $\tau\mapsto-f_{*}^{\prime}(\max\{(g({\bm{x}},{\bm{y}}^{\star})-\tau,f^{\prime}(0)\})$ is increasing on $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$ and therefore $\tau\mapsto h^{\prime}(\tau,{\bm{x}})=1-\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})f_{*}^{\prime}(\max\{g({\bm{x}},{\bm{y}})-\tau,f^{\prime}(0)\})$ is increasing on $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$.
Overall $h$ is well defined and strictly convex on $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$ such that $h^{\prime}(\tau_{\min}({\bm{x}}))\leq 0$ and $h^{\prime}(\tau_{\max}({\bm{x}}))\geq 0$. Hence, since $h$ is convex, we have
@@
and the unique minimizer $\tau^{\star}({\bm{x}})$ can then be found by solving the first order optimality condition $h^{\prime}(\tau,{\bm{x}})=0$ in $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$. We can therefore define a function $\tau^{\star}$ mapping each $\bm{x}$} to the unique root of $h^{\prime}(\tau,{\bm{x}})$ in $\tau_{\min}({\bm{x}}),\tau_{\max}({\bm{x}})]$.